OpenStudy (anonymous):

Translate the sentence into an equation. Six less than a number is 14. A. 6 – n = 14 B. 6n = 14 C. n – 6 = 14 D. n = 14 – 6

Mehek (mehek14):

Hey! What do you think it is?

OpenStudy (anonymous):

i think it is A

Mehek (mehek14):

think again 6 is less than n > 6 meaning n is bigger

OpenStudy (anonymous):

c right?

Mehek (mehek14):

yes

OpenStudy (anonymous):

i have 5 more to do would you like to help?

Mehek (mehek14):

sure

OpenStudy (anonymous):

Translate the sentence into an equation. The sum of 3 times a number and 7 is 19. A. 3(n + 7) = 19 B. (3 + n ) + 7 = 19 C. 7n + 3 = 19 D. 3n + 7 = 19

Mehek (mehek14):

what do you think?

OpenStudy (anonymous):

this one is a hard one for me

OpenStudy (anonymous):

i think it is D

OpenStudy (bohotness):

do you still nneed help?

OpenStudy (anonymous):

the more the marryer

Mehek (mehek14):

no 3 times a number and 7 so n+7 come first

Mehek (mehek14):

so now what do you think?

OpenStudy (anonymous):

i think it is A?

Mehek (mehek14):

yes you are correct :]

OpenStudy (anonymous):

Translate the sentence into an equation. The sum of 3 times a number and 7 is 19. A. 3(n + 7) = 19 B. (3 + n ) + 7 = 19 C. 7n + 3 = 19 D. 3n + 7 = 19

OpenStudy (anonymous):

wrong one

OpenStudy (anonymous):

Translate the sentence into an equation. Four times the difference of a number and 5 is 24. A. 4 • (c – 5) = 24 B. 4c – 5 = 24 C. c – 4 • 5 = 24 D. 5 • (c – 4) = 24

Mehek (mehek14):

a number and 5 come first so which one do you think it is?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

hello

OpenStudy (haifurriesftw):

No, it is A. 4*(c-5) = 24 4 times ([a number] minus 5) equals 24

OpenStudy (anonymous):

i have 2 more left

OpenStudy (anonymous):

Choose the equation that models the word problem. A football team won 33 games in the first three seasons. The team played a total of 48 games over the three seasons. How many games did they lose overall? Let g represent the number of games that the team did not win in the first three seasons. A. g – 33 = 48 B. g – 48 = 33 C. g + 33 = 48 D. 33g = 48

OpenStudy (haifurriesftw):

[number of games lost] + [number of games won] = [total number of games] you get it now?

OpenStudy (anonymous):

okay

OpenStudy (haifurriesftw):

which one do you think it is?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

last one

OpenStudy (haifurriesftw):

was making sure but yesh good, you are correct

OpenStudy (anonymous):

the last one is ezay would you like to see

OpenStudy (haifurriesftw):

i suppose but i dunno if i have time for that

OpenStudy (anonymous):

okay no problem

OpenStudy (haifurriesftw):

well post it if you like